The following data are available for two divisions of Solomons Company North Division South Division Division operating profiReg A1 Reg A2 Reg 1 Reg B2 Reg Reg C2 of Solomon measures performance using ROI, which division had the better performance? NComplete this question by entering your answers in the tabs below. Reg A1 Req AZ Req B1 Reg B2 Reg C Reg C2 Calculate the EVAComplete this question by entering your answers in the tabs below. RAL ROGAR B1 1 RB2 RCI Rec2 If Solomons measures performanComplete this question by entering your answers in the tabs below. Reg A1 Reg A2 Reg 1 Req B2 Reg CI Reg C2 Would your evaluaComplete this question by entering your answers in the tabs below Regal Reg A2 Req B1 ReqB2 Reg Ca Reg2 Would your evaluation
Answer #1
Operating Profit     6,355,000     41,730,000
Investment 31,000,000 321,000,000
a-1 ROI = Operating Income/Operating Investment 20.50% 13.00%
a-2 North Division perfomed Better
b-1 EVA = Operating Profit - Investment*Cost of Capital     3,565,000     12,840,000
b-2 South Division since higher EVA
c-1 NO, since Cost of capital does not affect ROI
EVA     1,085,000 (12,840,000)
Yes, since North Division is better now

в с North Divisio South Divisio 6355000 41730000 31000000321000000 1 2 Operating Profit 3 Investment 4. 5 a-1 ROI = Operating